OpenStudy (anonymous):

What is the length of measure EF and the measure of angle D. 10;37 17;53 10;53 14;37

OpenStudy (anonymous):

Here is the attachment

OpenStudy (anonymous):

pythagorean theorem

OpenStudy (anonymous):

17^2=EF^2 + 14^2

OpenStudy (anonymous):

the sum of angles in a triangle are 180 so 90+53+D=180

OpenStudy (anonymous):

Wait. I am confused. Which answer is it?

OpenStudy (anonymous):

i didnt give you the answer i give you the equation that needs only 1 step more to get the answer

OpenStudy (anonymous):

Are you able to tell me that step or will that give away the answer?

OpenStudy (anonymous):

it will give away the answer...

OpenStudy (anonymous):

17;53

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

I guessed. I have absolutely no clue what we are doing. I am not stupid! I just suck at math! Like, i really suck!

OpenStudy (anonymous):

ok D=37 but try to find EF

OpenStudy (anonymous):

34

OpenStudy (anonymous):

why 34?

OpenStudy (anonymous):

Because 17 x 2 =4 but that is completely wrong so forget that i said that

OpenStudy (anonymous):

17^2=EF^2 + 14^2

OpenStudy (anonymous):